Hello! I was wondering if I can have some help on this question here
Salim wants to download a certain number of his favourite tunes using some money he has saved doing odd jobs. At downloads for U, they charge $1.00 per download. He would need to save $17.00 more than he already has in order to download all he wants.
At a competing site, Music bytes, they only charge $0.80 per download. he would have $5.00 left from his savings of he downloads all he wants. How many downloads does Salim want?
I’m pretty sure this is a problem where you would use y=mx+b such as 17+1.00x=y but I’m not entirely sure, some direction would be nice. Thank you for your time!

Found 2 solutions by greenestamps, addingup:
Answer by greenestamps(13200)



Let's look at the equation you show: y = 1.00x+17.

The 1.00 is the cost per download at the first site; so 1.00x is the cost of downloading x tunes from that site. Then, since the amount of money Salim would need to download all his tunes from that site is $17 more than what he has, 1.00x+17 is the amount of money he would need to download all his tunes from that site.

So the equation you show is a good one for showing an expression for the amount of money Salim has.

Now do the same thing for the other site. At $0.80 per download from the other site, the amount of money he has is $5.00 MORE than what he needs. Use those numbers to make another equation in the form y = mx+b for the amount of money he has.

Then you will have two different expressions for the amount of money Salim has. Write and solve the equation that says those two expressions are equal to each other.

Answer by addingup(3677)

U = 1.00 per download. He would need 17 more
M = 0.80 per download. He would have 5 left over
1.00x - 17.00 = 0.80x + 5
0.2x = 22
x = 22/0.2 = 110
Salim wants 110 downloads.